# RegD: Hierarchical Embeddings via Distances over Geometric Regions

### v1

@@ 6.2 Inference over Ontologies
We utilize three normalized biomedical ontologies: GALEN (Rector et al., 1996), Gene Ontology (GO) (Ashburner et al., 2000), and Anatomy (Uberon) (Mungall et al., 2012). As in (Jackermeier et al., 2024), we use the entire ontology for training, and the complete set of inferred class subsumptions for testing. Those subsumptions can be regarded as partial order pairs $u\prec v$. Evaluation is performed using 1,000 subsumptions randomly sampled from the test set. Similar to inference over DAG, negative samples are generated by randomly replacing the child of each positive pair 10 times.
Baselines
We focus on the most representative ontology embedding methods: ELBE (Peng et al., 2022) and ELEM (Kulmanov et al., 2019), as well as their enhanced versions incorporating RegD or $\tau$Box. Other hierarchy embedding methods are excluded from our tests due to their incompatibility with ontology embeddings. For example, OE and EntailmentCone utilizes cones as embedding objects, which cannot be directly integrated with ELBE or ELEM for ontology tasks. Details of the integration are provided in Appendix F.2.
Results
The results are summarized in Tables 3. We can see that RegD yields consistent improvements across all ontologies for inferencing tasks. Specifically, it gains—an F1 score increase of more than 45% with ELBE method on the ANATOMY ontology. Conversely, the $\tau$Box plugin consistently reduces performance across nearly all test cases, underscoring its limited applicability to tasks involving more than hierarchies.
Table 4: F1 score (%) for the prediction task on ontologies.
Dataset Dim GALEN GO ANATOMY
ELBE d=5 21.0 32.8 25.1
    + $\tau$Box 18.4 24.0 25.6
    + RegD 20.6 37.1 41.4
ELBE d=10 24.2 37.9 25.5
    + $\tau$Box 19.5 29.3 22.9
    + RegD 21.0 44.1 45.3
ELEM d=5 16.7 54.5 23.1
    + RegD 16.8 60.3 21.7
ELEM d=10 16.6 54.2 26.0
    + RegD 18.0 61.4 21.9
@@ 6.3 Link Prediction over Ontologies
We use the same baselines and datasets as described in Section 6.2. However, in this prediction task, we partition the original ontologies directly into 80% for training, 10% for validation, and 10% for testing as in (Jackermeier et al., 2024). For the link prediction task, we focus on specific parts of the validation and testing sets, represented as $\exists r.B\prec?A$, where $A$ and $B$ are concept names and $r$ is a role. This setup is equivalent to link prediction tasks (?$A,r,B)$ in knowledge graphs if we regard $A$, $B$, and $r$ as the head entity, tail entity, and relation, respectively.
Results
Table 3 summarizes the results. RegD shows mixed results: while it generally improves performance, it decreases scores on the GALEN ontology and ANATOMY ontology with ELEM. This degradation likely occurs in challenging prediction cases where all methods perform poorly. Nevertheless, RegD achieves significant improvements in other cases, notably increasing F1 score by 77.6% with ELBE on the ANATOMY ontology. In contrast, the $\tau$Box plugin consistently reduces performance across almost all test cases.
@@ 7 Conclusion
We introduced a framework RegD for low-dimensional embeddings of hierarchies, leveraging two distance metrics between regions. Our method, applicable to regions in the Euclidean space, demonstrates versatility and has the potential for a wide range of tasks involving data beyond hierarchies. Additionally, we showed that our approach achieves comparable embedding performance to hyperbolic methods while being significantly simpler to implement.
For future work, we aim to extend the applicability of our methods to diverse geometric regions, tailoring their geometric properties to the specific requirements of different tasks. Moreover, we plan to explore additional applications beyond hierarchical structures, including integrating our framework with language models, as in (He et al., 2024b).
@@ A.1 Theorem 1
    Assume $\textit{reg}_{1}=\textit{ball}(\mathbf{c}_{1},r_{1})$ and $\textit{reg}_{2}=\textit{ball}(\mathbf{c}_{2},r_{2})$. For any ball $\textit{ball}(\mathbf{c}^{\prime},r^{\prime})\subseteq\textit{ball}(\mathbf{c}_{2},r_{2})$ with $r^{\prime}\leq 0.5\cdot r_{1}$, we must have:
    $\sum_{i}|\mathbf{c}_{i}-\mathbf{c}^{\prime}_{i}|^{p}+|r-r^{\prime}|^{p}>(0.5\cdot r_{1})^{p}.$
@@
    $d_{\text{dep}}(\textit{reg}_{1},\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))=\frac{\sum_{i}|\mathbf{c}_{i}-\mathbf{c}^{\prime}_{i}|^{p}+|r-r^{\prime}|^{p}}{f(\textit{ball})f(\textit{ball}^{\prime})}>\frac{\left(\frac{1}{2}r_{1}\right)^{p}}{f(\textit{ball})f(\textit{ball}^{\prime})}.$
@@
    $f(\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))<\epsilon:=\frac{f(\textit{ball})f(\textit{ball}^{\prime})[d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})+N]}{\left(\frac{1}{2}r_{1}\right)^{p}},$
    we can guarantee that $d_{\text{dep}}(\textit{reg}_{1},\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))>d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})+N$.
@@
    Therefore, there exists a $\delta>0$ such that when $r<\delta$, we have $f(\textit{reg})\leq\epsilon$ for $\textit{reg}=\textit{ball}(\mathbf{c},r)$. In conclusion, $r_{0}=\min\{\delta,0.5\cdot r_{1}\}$ satisfies the required condition.
@@
    Assume $\textit{reg}=\textit{ball}(\mathbf{c},r)$. For any $n,M>0$, we can select $n$ distinct vectors $\mathbf{c}_{1},\ldots,\mathbf{c}_{n}\in\textit{ball}(\mathbf{c},0.5\cdot r)$ such that:
@@
    Similarly to item 1, we can choose $\delta$ small enough such that for any ball $\textit{ball}(\mathbf{c}_{i},r_{i})$ with $r_{i}<\delta$, we have:
@@
The proof for the case of boxes follows the same reasoning, with the radius $r$ replaced by the offset $\mathbf{o}$ or its norm $||\mathbf{o}||$. ∎
@@ A.2 Proposition 1
Recall that when the curvature is $-1$, the distance in the hyperbolic space (half-plane model) is given by:
$d_{\mathbb{H}}(\mathbf{x},\mathbf{y})=\operatorname{arcosh}\left(1+\frac{\|\mathbf{x}-\mathbf{y}\|^{2}}{2\mathbf{x}_{n}\mathbf{y}_{n}}\right).$
The distance induced by the function $F$ is of the form:
$d^{\#}_{\mathcal{B}^{n}}(\textit{ball}(\mathbf{c},r),\textit{ball}^{\prime}(\mathbf{c}^{\prime}.r^{\prime}))=\operatorname{arcosh}\left(1+\frac{\|\mathbf{c}-\mathbf{c}^{\prime}\|^{2}+(r-r^{\prime})^{2}}{2rr^{\prime}}\right).$
This coincides with the depth distance in Example 1 when $p=2$, $g(x)=\operatorname{arcosh}(x+1)$, and $h(\textit{ball})=\sqrt{2}$. This completes the proof. ∎
@@ A.4 Proposition 3
Appendix B Hyperbolic Boundary Distance for Balls
@@
$\textit{Cone}(\textit{ball}(\mathbf{c},r),k)=\{(1-t)\mathbf{x}+t(\mathbf{c}+k\textbf{e}_{n+1}):\mathbf{x}\in\textit{ball}(\mathbf{c},r),t\in[0,1]\},\quad\text{where }\textbf{e}_{n+1}=(0,\ldots,0,1)\in\mathbb{R}^{n+1}.$
The mapping $G$ is defined by:
@@
[…]
@@
The boundary distance between two cones in the Poincaré half-plane model (as shown on the right of Figure 4) can be defined as:
@@
where $\chi=-1$ if $\textit{Cone}_{2}\subseteq\textit{Cone}_{1}$, and $\chi=1$ otherwise. $\partial(\textit{Cone}_{1})$ denotes the boundary of $\textit{Cone}_{1}$, and $\text{Apex}(\textit{Cone}_{1})$ is the apex of $\textit{Cone}_{1}$.
Assuming the curvature is fixed at $-1$ (i.e., $k=1$), we specialize Theorem 4.2 from Yu et al. (2024) by imposing the condition $\sinh(\sqrt{k}\,r)=1$, from which we derive:
@@
This distance can be extended back to the context of balls, allowing the definition of a boundary distance between balls as:
@@
Appendix C Poincare Ball Models
@@
$d_{\mathbb{H}}(x,y)=\frac{1}{\sqrt{k}}\,\text{arcosh}\left(1+\frac{2k\|x-y\|^{2}}{(1-k\|x\|^{2})(1-k\|y\|^{2})}\right).$
Appendix D $\delta$-Hyperbolicity
In the study of metric spaces, a space is said to be $\delta$-hyperbolic if it satisfies a particular property related to the triangle inequality. The concept of $\delta$-hyperbolicity was introduced by M. Gromov (Gromov, 1987) to capture the notion of negative curvature in a metric space.
Definition D.1 ($\delta$-Hyperbolicity).
Let $(X,d)$ be a metric space, where $d$ is a distance function between points in $x$. For any three points $x,y,z\in X$, the Gromov product is defined by:
$(y\cdot z)_{x}=\frac{1}{2}(d(x,y)+d(x,z)-d(y,z))$
The space is called $\delta$-hyperbolic if there exists a constant $\delta\geq 0$ such that for all points $x,y,z,w\in X$:
$(y\cdot z)_{x}\geq\min\{(y\cdot w)_{x},(w\cdot z)_{x}\}-\delta$
Generally, a $\delta$-hyperbolic space approximates a hyperbolic geometry when $\delta$ is small, and it approximates Euclidean geometry as $\delta$ grows large. Moreover, $\delta$-hyperbolicity can also be applied to discrete spaces such as graphs, with the distance $d(x,y)$ is the shortest path distance between two vertices $x$ and $y$. In this context, $\delta$-hyperbolicity helps capture the ”thickness” of the graph’s structure and can be used to analyze its intrinsic geometry.
Approximate Computing Process of $\delta$-Hyperbolicity
$\delta$-Hyperbolicity can be computed approximately by randomly sampling a subset of quadruples. Formally, by sampling $N$ quadruples $(x,y,z,w)$ from a set $S$, the approximate value of $\delta$-hyperbolicity can be computed as
$\delta_{\text{app}}=\max\left\{\min\{(y\cdot w)_{x},(w\cdot z)_{x}\}-(y\cdot z)_{x}\mid(x,y,z,w)\in S\right\}.$
To provide a more robust estimate, we also compute the mean value of the $\delta$-hyperbolicity across all samples. This allows us to capture both the maximal influence of local behavior and the overall average property of the space. For example, in Table 2, we approximate the $\delta$-hyperbolicity by sampling 50,000 quadruples and report both the maximum and the mean of the sampling results to offer a comprehensive understanding of the space’s geometric characteristics.

### v2

@@ 4.2 Inference and Link Prediction over Ontologies
Benchmark We utilize three normalized biomedical ontologies: GALEN Rector et al. (1996), Gene Ontology (GO) Ashburner et al. (2000), and Anatomy (Uberon) Mungall et al. (2012). As in Jackermeier et al. (2024), we use the entire ontology for training, and the complete set of inferred class subsumptions for testing. Those subsumptions can be regarded as partial order pairs $u\prec v$. Evaluation is performed using 1,000 subsumptions randomly sampled from the test set. Similar to inference over DAG, negative samples are generated by randomly replacing the child of each positive pair 10 times.
[…]
Table 3: F1 score (%) for the inference task.
Method GALEN GO ANATOMY
d=5 d=10 d=5 d=10 d=5 d=10
ELBE 20.7 21.2 36.9 42.4 43.1 43.0
[…]
+ RegD 25.2 25.8 50.0 50.5 58.7 62.5
ELEM 16.9 17.3 23.5 27.4 34.6 38.7
+ RegD 19.2 18.8 36.4 40.0 52.5 55.5
Table 4: F1 score (%) for the prediction task.
Method GALEN GO ANATOMY
d=5 d=10 d=5 d=10 d=5 d=10
ELBE 21.0 24.2 32.8 37.9 25.1 25.5
[…]
+ RegD 20.6 21.0 37.1 44.1 41.4 45.3
ELEM 16.7 16.6 54.5 54.2 23.1 26.0
+ RegD 16.8 18.0 60.3 61.4 21.7 21.9
@@ 4.3 Link Prediction over Ontologies
We use the same baselines and datasets as described in Section 4.2. However, in this prediction task, we partition the original ontologies directly into 80% for training, 10% for validation, and 10% for testing as in Jackermeier et al. (2024). For the link prediction task, we focus on specific parts of the validation and testing sets, represented as $\exists r.B\prec?A$, where $A$ and $B$ are concept names and $r$ is a role. This setup is equivalent to link prediction tasks $(?A,r,B)$ in knowledge graphs if we regard $A$, $B$, and $r$ as the head entity, tail entity, and relation, respectively.
[…]
@@ 5 Conclusion
We introduced a framework RegD for low-dimensional embeddings of hierarchies, leveraging two dissimilarity metrics between regions. Our method, applicable to regions in the Euclidean space, demonstrates versatility and has the potential for a wide range of tasks involving data beyond hierarchies. Additionally, we showed that our approach achieves comparable embedding performance to hyperbolic methods while being significantly simpler to implement.
For future work, we are interested in integrating our method with other approaches based on hyperbolic spaces. For instance, Hyperbolic Neural Networks Ganea et al. (2018a) currently depend on intricate manifold operations, such as parallel transport, to perform matrix multiplications and vector additions. In contrast, our framework enables these computations to be carried out more directly through the parameter space, potentially simplifying the implementation. Another promising direction for future work is to investigate replacing existing hyperbolic embedding components in related methods with our approach, including recent studies that couple large language models with hyperbolic embeddings to learn semantic hierarchies He et al. (2024b). Also, we are interested in generalizing our framework to support a wider variety of region types—beyond the balls and boxes considered here—so that it can be better adapted to the requirements of diverse downstream tasks.
Reproducibility statement
All code and data are publicly available at https://anonymous.4open.science/r/RegD-F4E3, along with clear instructions on environment configuration and hyperparameter settings to enable full reproduction of our results.
@@ A.1 Theorem 1
    Let $\textit{reg}_{1}=\textit{ball}(\mathbf{c}_{1},r_{1})$ and $\textit{reg}_{2}=\textit{ball}(\mathbf{c}_{2},r_{2})$. For any ball $\textit{ball}(\mathbf{c}^{\prime},r^{\prime})\subseteq\textit{ball}(\mathbf{c}_{2},r_{2})$ with $r^{\prime}\leq 0.5\cdot r_{1}$, we must have:
    $\mathbf{c}_{1}-\mathbf{c}^{\prime}||^{p}_{p}+|r_{1}-r^{\prime}|^{p}>(0.5\cdot r_{1})^{p}.$
@@
[…]
@@
    $f(\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))<\epsilon:=\frac{\left(0.5\cdot r_{1}\right)^{p}}{f(\textit{reg}_{1})[d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})+\Delta]},$
    we can guarantee that $d_{\text{dep}}(\textit{reg}_{1},\textit{ball}(\mathbf{c}^{\prime},r^{\prime}))>d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})+\Delta$.
@@
    Therefore, there exists a $\delta>0$ such that when $r<\delta$, we have $f(\textit{reg})\leq\epsilon$ for any region $\textit{reg}=\textit{ball}(\mathbf{c},r)$. In conclusion, $r_{0}=\min\{\delta,0.5\cdot r_{1}\}$ satisfies the required condition.
@@
    Let $\textit{reg}=\textit{ball}(\mathbf{c},r)$. For any $n,M>0$, we can select $n$ distinct vectors $\mathbf{c}_{1},\ldots,\mathbf{c}_{n}\in\textit{ball}(\mathbf{c},0.5\cdot r)$ such that:
@@
    Similarly to item 1, we can choose $r_{i}$ small enough such that for any ball $\textit{ball}(\mathbf{c}_{i},r_{i})$, we have:
@@
The proof for the case of boxes follows the same reasoning, with the radius $r$ replaced by the offset $\mathbf{o}$ or its norm $\mathbf{o}||$. ∎
@@ A.2 Proposition 1
Recall that when the curvature is 1-1, the distance in the hyperbolic space (half-plane model) is given by:
$d_{k}(\mathbf{x},\mathbf{y})=\operatorname{arcosh}\left(1+\frac{\|\mathbf{x}-\mathbf{y}\|^{2}_{2}}{2\mathbf{x}_{n}\mathbf{y}_{n}}\right).$
Then, the distance induced by the function $F$ is of the form:
$d^{\#}_{\mathcal{B}^{n}}(\textit{ball}(\mathbf{c},r),\textit{ball}^{\prime}(\mathbf{c}^{\prime}.r^{\prime}))=\operatorname{arcosh}\left(1+\frac{\|\mathbf{c}-\mathbf{c}^{\prime}\|^{2}_{2}+(r-r^{\prime})^{2}}{2rr^{\prime}}\right).$
This coincides with the depth dissimilarity in Example 1 when $p=2$, $g(x)=\operatorname{arcosh}(x+1)$, and $f(\textit{ball}(\mathbf{c},r))=\sqrt{2}\,r$. This completes the proof. ∎
@@ A.4 Proposition 3
A.5 General Parameterized Regions
A set of parameterized regions over Euclidean Space $\mathbb{R}^{n}$ can be defined as all regions reg of the following form:
$\textit{reg}=\{\mathbf{x}\in\mathbb{R}^{n}:f_{i}(\mathbf{x},\theta)\leq 0,\quad\text{for }i=1,\ldots,n\},$
where $f_{1},\ldots,f_{n}$ are given (differentiable) functions from the given space to $\mathbb{R}$ and $\theta\in\mathbb{R}^{m}$ is the $m$-dimensional parameter.
Notably, the results established in Theorem 1 naturally generalize to other parameterized regions, with proofs following analogous arguments. The result for the general case is of the following form.
Theorem 2.
Consider the parameterized region space $\mathcal{R}$ equipped with the depth dissimilarity as defined in Definition 1. The following properties hold:
  1. 1.
    For any $\textit{reg}_{1},\textit{reg}_{2}\in\mathcal{R}$ and any $\Delta>0$,for any subset $\textit{reg}^{\prime}\subseteq\textit{reg}_{2}$ $\textit{reg}_{1}$ sufficiently small, we have
    $d_{\text{dep}}(\textit{reg}_{1},\textit{reg}^{\prime})>d_{\text{dep}}(\textit{reg}_{1},\textit{reg}_{2})+\Delta.$
  2. 2.
    For any $\textit{reg}\in\mathcal{R}$, any integer $n$, and any $M>0$, there exist subsets $\textit{reg}_{1},\ldots,\textit{reg}_{n}\subseteq\textit{reg}$ such that for any distinct $i,j\in\{1,\ldots,n\}$,
    $d_{\text{dep}}(\textit{reg}_{i},\textit{reg}_{j})>M.$
Outline of proof.
We sketch the main ideas, which closely follow the argument in the proof of Theorem 1.
For the first statement, by selecting a sufficiently small subset $\textit{reg}^{\prime}\subseteq\textit{reg}_{2}$, we can ensure that
$\|\textit{P}(\textit{reg})-\textit{P}(\textit{reg}_{1})\|>\delta$
for some $\delta>0$. Additionally, since $f(\textit{reg}^{\prime})$ can be made arbitrarily small as $\textit{reg}^{\prime}\to\emptyset$ (i.e., $\lim_{\textit{reg}\to\emptyset}f(\textit{reg})=0$), we can make the numerator of the depth dissimilarity arbitrarily small while keeping the denominator bounded below by a positive constant. Consequently, the depth dissimilarity $d_{\text{dep}}(\textit{reg}_{1},\textit{reg}^{\prime})$ can be made arbitrarily large.
For the second statement, a similar argument applies: by choosing the subsets $\textit{reg}_{1},\ldots,\textit{reg}_{n}$ to be sufficiently small and mutually separated within reg, we can ensure that each $f(\textit{reg}_{i})$ is small enough so that the depth dissimilarity between any pair $(\textit{reg}_{i},\textit{reg}_{j})$ exceeds $M$.
For Proposition 1, the mapping $F$ above can be generalized to arbitrary regions. For example, we could extend $F$ by replacing the center point and radius of balls with the center of gravity and diameters of any region. However, in this case, the function $F$ might be non-injective, as distinct regions may share the same center of gravity and diameters. Therefore, the resulting region space serves more as an extension rather than an equivalence of hyperbolic space.
Appendix B Hyperbolic Boundary Dissimilarity for Balls
@@
$\textit{Cone}(\textit{ball}(\mathbf{c},r),k)=\{(1-t)\mathbf{x}+t(\mathbf{c}+k\textbf{e}_{n+1}):\mathbf{x}\in\textit{ball}(\mathbf{c},r),t\in[0,1]\},$
where $\textbf{e}_{n+1}=(0,\ldots,0,1)\in\mathbb{R}^{n+1}$. The mapping $G$ is defined by:
@@
$\mathbb{R}$1-111xxyy1-11111ball on $\mathbb{R}$cone over upper half-space of $\mathbb{R}^{2}$
@@
The boundary dissimilarity between two cones in the Poincaré half-plane model (as shown on the right of Figure 4) can be defined as:
@@
where $\chi=-1$ if $\textit{Cone}_{2}\subseteq\textit{Cone}_{1}$, and $\chi=1$ otherwise. $\partial(\textit{Cone}_{1})$ denotes the boundary of $\textit{Cone}_{1}$, and $\text{Apex}(\textit{Cone}_{1})$ is the apex of $\textit{Cone}_{1}$, which is a single point.
Assuming the curvature is fixed at 1-1 (i.e., $k=1$), we specialize Theorem 4.2 from Yu et al. (2024) by imposing the condition $\sinh(\sqrt{k}\,r)=1$, from which we derive:
@@
This distance can be extended back to the context of balls, allowing the definition of a boundary dissimilarity between balls as:
@@
Appendix C Poincaré Ball Models
@@
$d_{k}(x,y)=\frac{1}{\sqrt{k}}\,\text{arcosh}\left(1+\frac{2k\|x-y\|^{2}}{(1-k\|x\|^{2})(1-k\|y\|^{2})}\right).$
Appendix D Experiments Setting
D.1 Hyperparameters
Following the setup in Yu et al. (2024), we evaluated our models in dimensions $5$ and $10$. The weighting coefficient $\lambda$ in Eq. 7 was selected from $\{0,0.5,1\}$, while the learning rate was tuned over $\{0.01,0.005,0.001,0.0005\}$. For the small dataset of mammals, which contains approximately 1000 nodes, we choose batch sizes from $\{32,64,128\}$. For the remaining datasets, we opt for batch sizes from $\{1024,2048,8192\}$. For the depth dissimilarity, we tested $p\in\{1,2\}$ as specified in Definition 2. The boundary dissimilarity margins in Eq. 4 were set to $\gamma_{1}=0.001$ and $\gamma_{2}=0$ for directed acyclic graph (DAG) embedding tasks. In ontology embedding experiments, both margins were set to zero.
For DAG experiments, as in Yu et al. (2024), we trained for 400 epochs on the mammal and noun datasets and 500 epochs on the MCG and Hearst datasets. For ontology embedding experiments, we trained for 5000 epochs, following the protocol in the previous work Jackermeier et al. (2024) of ontology embedding.
All experiments were conducted on a system equipped with an AMD Ryzen Threadripper PRO 7965WX 24-core processor, 128GB of RAM, and an NVIDIA A6000 GPU with 48GB of VRAM, running Ubuntu 24.04.
D.2 Ablation Studies
The results of the ablation studies are summarized in Table 5. Note that, when $\lambda=0$, the depth dissimilarity component is absent from the energy function, and therefore, the choice of $p$, as a hyperparameter of depth dissimilarity, is not applicable.
From the results, it is evident that our boundary dissimilarity generally outperforms the specific boundary dissimilaritys for boxes and balls (i.e., $d_{\textit{vol}}$ and $d_{\text{bd}}^{\text{cone}}$) in most cases. Furthermore, when using boxes as the embedding object, setting $p=1$ yields better results than $p=2$. However, for balls, $p=2$ achieves superior performance.
For the mammal and noun datasets, using balls as the embedding object and incorporating the depth dissimilarity (i.e., $\lambda>0$) in the energy function results in better performance. Conversely, for the MCG and Hearst datasets, the best performance is typically achieved when using boxes as the embedding region and setting $\lambda=0$, effectively excluding the hyperbolic components. This observation aligns with our analysis in the main text.
It is worth noting, as shown in the seventh row of Table 5, that even under a small configuration range (i.e., with box and $\lambda\in{0.5,1}$), our method outperforms existing approaches and exhibits only minimal performance gaps compared to the best results obtained with other configurations.
Options Mammal Noun MCG Hearst
region $\lambda$ p boundary dissimilarity d=2 d=5 d=5 d=10 d=5 d=10 d=5 d=10
Box 0 - $d_{\text{bd}}$ 39.4 42.8 34.6 34.6 49.9 56.6 41.6 44.7
$d_{\textit{vol}}$ 31.7 38.1 30.8 31.9 48.6 58.5 42.8 49.6
>0>0 1 $d_{\text{bd}}$ 64.9 71.6 53.8 51.3 43.7 45.0 35.6 38.8
$d_{\textit{vol}}$ 45.6 54.3 30.4 33.4 30.9 39.9 32.6 38.8
2 $d_{\text{bd}}$ 56.1 62.2 50.3 47.4 50.7 57.4 40.6 44.6
$d_{\textit{vol}}$ 39.7 47.9 30.6 31.6 48.5 56.7 37.8 43.8
Ball 0 - $d_{\text{bd}}$ 42.8 55.5 39.6 42.5 43.9 46.7 37.7 37.1
$d_{\text{bd}}^{\text{cone}}$ 41.2 49.7 36.4 37.7 40.9 43.7 36.0 36.7
>0>0 1 $d_{\text{bd}}$ 58.2 69.2 51.7 54.4 40.9 45.9 33.6 35.8
$d_{\text{bd}}^{\text{cone}}$ 57.9 65.3 52.3 54.7 36.8 41.7 31.6 34.6
2 $d_{\text{bd}}$ 62.5 71.8 58.4 59.1 41.7 45.8 34.8 37.7
$d_{\text{bd}}^{\text{cone}}$ 62.7 67.7 54.3 55.1 40.7 45.7 34.6 36.7
D.3 Other Results
Including non-basic edges in training
Table 6: F1 score (%) on mammal sub-graph with best numbers bolded. Results with * are coming from Yu et al. (2024). For ShadowCone, only result of Umbral-half-space is reported as in Yu et al. (2024).
Non-basic-edge Dimension = 2 Dimension = 5
Percentage 0% 10% 25% 50% 90% 0% 10% 25% 50% 90%
GBC-box 23.4 25.0 23.7 43.1 48.2 35.8 60.1 66.8 83.8 97.6
VBC-box 20.1 26.1 31.0 33.3 34.7 30.9 43.1 58.6 74.9 69.3
[…]
EntailmentCone 54.4 61.0 71.0 66.5 73.1 56.3 81.0 84.1 83.6 82.9
ShadowCone 57.7 73.7 77.4 80.3 79.0 69.4 81.1 83.7 88.5 91.8
Ours (box) 64.9 74.4 75.8 78.3 82.8 71.6 77.8 83.2 87.6 87.0
Ours (ball) 62.7 68.3 77.2 84.1 88.6 71.8 74.5 84.4 88.4 90.2
Repeatness
The main experiments presented in this paper were conducted using a fixed random seed, following established practices in previous works Yu et al. (2024); Ganea et al. (2018b). To assess the repeatability and robustness of our results, we conducted 10 independent experimental runs on each of the four datasets. The average performance and standard deviation are reported in Table 7, while the outcomes of individual runs are summarized in Figure 5. These results demonstrate that our method exhibits considerable stability, particularly on the MCG and Hearst datasets. Compared to the main results obtained with a fixed random seed, the overall average performance remains very close, with only a few exceptions:
  1. 1.
    In certain cases—such as using balls on the Mammal dataset with $d=2$ and on the Noun dataset with d=5d=5—performance is slightly lower, though still superior to that of other baseline methods.
  2. 2.
    In some settings, performance is even improved on average, for example, when using balls on the Noun dataset with $d=10$.
These results indicate that the proposed method is robust across different random seeds, with consistent performance across most settings.
Figure 5: All F1 scores of 10 random runs on Mammal, WordNet noun, MCG, and Hearst.
Table 7: Average and standard deviation (Std) of 10 random runs of F1 score (%) on Mammal, WordNet noun, MCG, and Hearst.
RegD Mammal Noun MCG Hearst
d=2 d=5 d=5 d=10 d=5 d=10 d=5 d=10
fixed seed box 64.9 71.6 53.8 51.3 50.7 58.5 42.8 49.6
ball 62.7 71.8 58.4 59.1 44.9 46.8 37.7 37.7
average box 63.00 69.65 52.98 49.33 50.49 57.71 42.70 48.72
[…]
average ball 56.70 66.76 51.63 59.31 44.09 46.35 37.24 37.11
[…]
Impact of $\lambda$
In Table 5, we evaluate the performance of $\lambda=0$, $0.5$, and $1$, and report results for both $\lambda=0$ and $\lambda>0$. To provide a more detailed analysis, Table 8 presents additional experiments with $\lambda$ ranging from 0 to $1$ in increments of $0.1$, using the mammal dataset (dim = 2, box embedding).
The results suggest that the impact of $\lambda$ may not follow a clear or consistent trend, and the optimal value of $\lambda$ can vary significantly across different settings. For instance, in this dataset, larger values of $\lambda$ generally lead to higher F1 scores compared to smaller ones. However, the relationship is not strictly monotonic—an increase in $\lambda$ does not always result in improved performance.
0 0.1 0.2 0.3 0.4 0.5 0.6 0.7 0.8 0.9 1
F1(avg) 42.70 50.92 57.42 56.07 57.03 64.43 62.46 62.83 66.56 62.80 65.30
(±std) (±7.07) (±5.44) (±5.46) (±4.87) (±5.62) (±4.61) (±3.23) (±2.11) (±2.33) (±2.47) (±2.08)
Table 8: F1 scores with standard deviations for different $\lambda$ values on the mammal dataset (dim=2, using box).